# Application of Whisper in Clinical Practice: the Post-Stroke Speech Assessment during a Naming Task thanks:

### v1

@@ Abstract
Detailed assessment of language impairment following stroke remains a cognitively complex and clinician-intensive task, limiting timely and scalable diagnosis. Automatic Speech Recognition (ASR) foundation models offer a promising pathway to augment human evaluation through intelligent systems, but their effectiveness in the context of speech and language impairment remains uncertain. In this study, we evaluate whether Whisper, a state-of-the-art ASR foundation model, can be applied to transcribe and analyze speech from patients with stroke during a commonly used picture-naming task. We assess both verbatim transcription accuracy and the model’s ability to support downstream prediction of language function, which has major implications for outcomes after stroke. Our results show that the baseline Whisper model performs poorly on single-word speech utterances. Nevertheless, fine-tuning Whisper significantly improves transcription accuracy (reducing Word Error Rate by 87.72% in healthy speech and 71.22% in speech from patients). Further, learned representations from the model enable accurate prediction of speech quality (average F1 Macro of 0.74 for healthy, 0.75 for patients). However, evaluations on an unseen (TORGO) dataset reveal limited generalizability, highlighting the inability of Whisper to perform zero-shot transcription of single-word utterances on out-of-domain clinical speech and emphasizing the need to adapt models to specific clinical populations. While challenges remain in cross-domain generalization, these findings highlight the potential of foundation models, when appropriately fine-tuned, to advance automated speech and language assessment and rehabilitation for stroke-related impairments.
Index Terms:
Automatic speech recognition, Clinical speech analysis, Language impairment detection
I Introduction
@@
Stroke is one of the leading causes of adult death and disability worldwide, with global incidence of 12.2 million per year [1]. One of the most debilitating and common consequences after stroke is aphasia, an acquired disorder of speech and language production and comprehension, with over 30% prevalence in patients with stroke[2, 3, 4]. Due to high heterogeneity of aphasic symptoms, diagnosis requires an in-depth assessment to uncover specific impairments and tailor speech therapy to individual patient’s needs. Computerized assessments have the advantage, compared to standard pen and paper tests, of being administered with minimal supervision, providing a cost-effective and accessible platform for speech and language evaluation of patients with stroke in varied settings [5]. However, these digital assessments often depend on manual speech transcription and evaluation by a trained clinician, which is time-consuming, prone to inconsistencies due to inter- and intra-rater variability and are difficult to implement on a large scale.
@@
Advancements in automatic speech recognition (ASR) methods can fill this gap and allow for automation of speech evaluation pipelines. However, whilst the latest ASR models perform exceedingly well on healthy speech data [8], several limitations have hindered the translation of ASR to pathological speech analysis. Firstly, currently available ASR models have been trained on healthy speech, limiting their generalizability to pathological speech. Secondly, the potential variability in speech impairments in aphasia necessitates large clinical speech databases for effective training of ASR models, which are currently relatively limited.
The extensive linguistic knowledge embedded in foundation ASR model, acquired through pretraining on large and diverse datasets, has the potential to be leveraged in a zero and few shot learning context to automate post-stroke speech evaluation pipelines [6, 7].
@@
In this study, we aimed to assess the performance of Whisper [8], a state-of-the-art transformer-based ASR model pretrained on 680,000 hours of multilingual data, on a stroke-specific speech database derived from a Naming task, commonly used in post-stroke speech assessments. We fine-tuned Whisper on speech from age-matched healthy older adults and patients with stroke and compared the verbatim transcription performance of the fine-tuned and the baseline Whisper models. Further, we evaluated whether Whisper-derived representations could be effectively applied to clinically-relevant downstream tasks, specifically speech impairment severity prediction. Lastly, we assessed the quality of these predictions by conducting a divergent and convergent validity analysis with clinical stroke outcomes.
II Related Work
@@
Current approaches for automated speech impairment analysis involve three key steps. First, speech-to-text ASR models are employed to generate speech transcriptions from audio recordings. Then, a feature extraction process is applied to the produced transcriptions and audio data. The extracted lexico-acoustic features are used in downstream analyses, such as classification and regression, to estimate the degree of speech impairment [9]. Different algorithms have been used to complete each one of these steps.
II-A Recurrent Neural Networks
@@
Previous studies have trained customized ASR models to transcribe aphasic speech, subsequently deriving acoustic and transcription-based features to estimate aphasia severity [11, 10]. Le and colleagues[11] utilized a Bidirectional Long-Short Memory-Recurrent Neural Network (BLSTM-RNN) model trained on acoustic features of aphasic speech derived from the AphasiaBank dataset [12] and frame-level senone and monophone labels, obtained through a Hidden Markov Model-Gaussian Mixture Model (HMM-GMM) system, to produce speech transcriptions. Similarly, Qin and colleagues[10] utilized a Time-Delay Neural Network combined with a Bidirectional Long-Short Term Memory (TDNN-BLSTM) model to transcribe Cantonese aphasic speech. The model was trained on acoustic features derived from healthy speech and HMM-GMM-derived triphone target labels [10]. The ASR transcriptions were used to extract features based on text alone [11], or in combination with audio features [10] to estimate aphasia severity by regression. While the above studies reached promising accuracy, the employed ASR systems had complex architectures, which required a multi-step training process, limiting global optimization [13]. Additionally, these models were based on variants of RNN architectures, which are constrained in modeling long-range dependencies compared to Transformers [14].
II-B Transformers
The advent of Transformers has revolutionized the field of ASR by implementing self-attention mechanisms, which capture temporal relationships across the entire input sequence simultaneously, thus overcoming the range limitations of RNN-based architectures and enabling more efficient encoding of contextual information [15, 16]. Multiple ASR transformer-based models are currently available for speech-to-text transcription including Transformer Transducer [17], wav2vec 2.0 [18], SpeechStew [19], Conformer [20] and Whisper [8]. However, while transformer-based models have demonstrated exceptional performance on healthy speech ASR [8], they often encounter challenges in adapting to aphasic speech recognition due to the high heterogeneity in the lexico-acoustic features of post-stroke speech.
@@ II-B Transformers
Recent studies have leveraged pretrained transformer-based ASR models, adapting them to enhance performance on aphasic speech. Within this framework[21], XLSR-53, a model built upon wav2vec 2.0 architecture, was fine-tuned to adapt it to aphasic speech transcription. XLSR-53 [22], pretrained on the extensive dataset, which contains 56,000 hours of unlabeled audio, uses convolutional neural network layers to generate latent speech representations [18]. These representations are then passed through a transformer network to encode contextualized speech features [18]. However, wav2vec 2.0 architecture is outperformed by other transformer-based foundation models such as Whisper [8, 23]. Whisper-derived representations have been successfully utilized in previous studies to classify dysfluencies in stuttering [24] and to estimate dysarthria severity [25]. Further, Whisper and XLSR-53 model transcripts have been successfully used for relevant speech feature extraction, later used for aphasia type classification, reaching average 90.6 F1 score [26]. Recent work also investigated hyperparameter tuning and fine-tuning of Whisper to better fit aphasic speech, reaching promising results of 38.5% and 21.93% WER on aphasic speech transcription, respectively. [27, 28]. However, the vast majority of previous work has focused on sentential speech, and the applications of Whisper to single-word utterances remain to be studied.
III Method
Figure 1: General study pipeline. Overview of the training and inference pipeline for verbatim transcription and accuracy prediction models. Transcription: Training – Whisper models were fine-tuned on speech data from 5 different datasets. A synthetic dataset of audio-transcription word pairs based on stimuli in the Naming task, SONIVA-Naming healthy dataset, collected from healthy age-matched participants performing the Naming task, SONIVA-Naming patient dataset, collected from patients with stroke performing the Naming task, a combined dataset of SONIVA-Naming healthy and patient data, and a combined dataset of all available data (synthetic, SONIVA-Naming healthy and SONIVA-Naming patient). All models were trained using cross-entropy loss to transcribe spoken words (e.g. “acorn”) from the Naming task input audio. Inference – Fine-tuned models (ft-syn, ft-h, ft-p, ft-hp and ft-all) were evaluated on 4 separate datasets – synthetic, SONIVA-Naming healthy, SONIVA-Naming patient, and an unseen testing dataset of dysarthric speech derived from the TORGO database. Accuracy: Training – accuracy prediction models were trained on SONIVA-Naming derived healthy and patient speech in a linear probing framework. The encoder layer of the models was frozen and the classification projection head was trained to predict accuracy scores using cross-entropy loss (e.g “2”). The encoder weights of the models were initialized at baseline or with the weights derived from the models trained for verbatim transcription of healthy or patient speech (ft-h or ft-p).
III-A Dataset
Speech data from age-matched controls and patients with stroke were derived from participants undergoing speech and cognitive testing as part of the the IC3 study [5]. Specifically the data was obtained from the Naming task, modified from the 30-item Boston Naming Task [29], where participants were instructed to name a depicted black and white line drawing picture with a single word. The speech recordings from this task, amongst other speech production tasks, have contributed to a larger speech database SONIVA (Speech Recognition Validation in Aphasia) [30]. The data used in this study will be referred to as SONIVA-Naming henceforth. The accuracy of the Naming task performance was evaluated across each trial using four accuracy metrics: semantic content (Semantic), fluency of articulation (Dysfluency), presence of self-corrections in the response (Self-correction) and phonological correctness (Phonology). Each trial was scored by trained raters on a 3-point scale: 0, 1, or 2, where 2 corresponds to the highest accuracy. The verbatim transcriptions of the audio files ranged from single word to sentences depending on the participants responses. The inter-rater reliability against a qualified speech therapist was high (Mean Intraclass Correlation Coefficient = 0.83$\pm$ 0.0046), and all ambiguous cases were resolved by consulting a speech therapist. In total, the SONIVA-Naming healthy dataset comprised 3960 Naming task trials from 132 healthy participants, while the SONIVA-Naming patient dataset consisted of 2609 trials from 87 patients (see Table I). There was no group difference for age (Mann-Whitney U statistic: $0.885542.50,p=$ 0.88). Patients had lower education ($\chi^{2}$=26.19, $p<0.001$) and a higher proportion of non-native English speakers ($\chi^{2}$=4.02, $p=0.04$). A noise-free, single word, baseline synthetic dataset was additionally generated using the Google Text-to-Speech library for Python [31]. Seven English accents were used to generate a total of 210 synthetic audio files across the 30 stimuli words available in the Naming task.
@@ III-A Dataset
An additional unseen testing dataset was used for assessment of model generalizability for single-word pathological speech transcription. This dataset was derived from the TORGO database [32]. The database includes speech recordings from individuals with dysarthria, a disorder caused by disruptions in the speech motor system, common in conditions such as cerebral palsy and amyotropic lateral sclerosis. The single-word recordings from the database were used to assess model transcription of short-form data. The TORGO dataset consisted of data from 7 participants (4 male and 3 female) and comprised a total of 1240 utterances.
TABLE I: Sample demographics of the SONIVA-Naming dataset
Demographics feature SONIVA-Naming healthy SONIVA-Naming patients
(N = 132) (N = 87)
Mean (Standard Deviation)
Age 61.6 (10.8) 61.8 (13.6)
Sex***
     Male:Female 52:77 62:25
     Missing data 3 0
English Language*
      Non-native 27 30
      Native 102 57
Education level***
     School 36 45
     Degree 43 37
     Post-graduate 37 3
     Missing data 16 2
${}^{\mathrm{**}}\;p<0.05$; ${}^{***}\,p<0.001$
III-B Data Pre-processing
@@
All audio data were resampled to 16 kHz. In the main dataset used for model training and evaluation (SONIVA-Naming), all audio recordings without at least one ground truth accuracy metric (Semantic, Dysfluency, Self-correction and Phonology) were excluded when fine-tuning models for the accuracy score prediction task ($n=23$ and $n=114$ for SONIVA-Naming healthy and SONIVA-Naming patient datasets, respectively). The data were divided into training, validation and test sets in the ratio of 7:1:2. The splitting was completed at the participant level such that all trials from the same participants were in the same partition. Due to the high class imbalance across the accuracy scores, the training sets for the accuracy prediction task were randomly downsampled to the minority class in each accuracy metric. This resulted in 144, 204, 90 and 21 audio files for the SONIVA-Naming healthy dataset for Semantic, Dysfluency, Self-correction and Phonology, respectively. In case of the SONIVA-Naming patient dataset, the final sample size was 198, 729, 108, and 171 trials for the same metrics respectively.
III-C Whisper Architecture
To complete the accuracy classification task, the same encoder architecture was used, but the decoder was replaced with a linear classifier projection head, consisting of two linear layers.
III-D Model Training
@@
Whisper was separately fine-tuned on the training set of synthetic, SONIVA-Naming healthy, and SONIVA-Naming patient data, as well as combined datasets of all participant data (SONIVA-Naming healthy and SONIVA-Naming patient) and all available data (synthetic, SONIVA-Naming healthy and SONIVA-Naming patient). Small and Medium Whisper models were used to determine the optimal model size. In total, ten fine-tuned models were obtained: fine-tuned on synthetic data (ft-syn), on SONIVA-Naming healthy data (ft-h), on SONIVA-Naming patient data (ft-p), on all participant data (ft-hp) and all available data (ft-all), each in Small and Medium sizes.
Models were trained for a maximum of 1000 steps, with a batch size of 16. A learning rate of $1\times 10^{-5}$ was used, with a linear learning rate scheduler incorporating 250 warm up steps. The model parameters were updated with AdamW optimiser using Cross-Entropy loss function (1).
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}\sum_{j=1}^{C}y_{ij}\log(\hat{y}_{ij}),$ (1)
where N is the total number of samples, C is the total number of classes, $y_{ij}$ is the true label and $\hat{y}_{ij}$ is the predicted probability for class $j$ for instance $i$.
Training was conducted on one NVIDIA RTX 6000 GPU. Models were evaluated on the validation set (batch size of 8) at 50 step intervals. The best model was chosen based on Word Error Rate (WER) on the validation set and was used for final evaluation on the test sets. The evaluation was completed on four different test sets, namely the synthetic, SONIVA-Naming healthy, and SONIVA-Naming patient as well as the unseen TORGO database.
@@
The setup was formulated as a multi-class classification task, where one of three possible accuracy scores (0, 1 or 2) were predicted for each given Naming task trial recording in the SONIVA-Naming healthy and patient datasets.
The encoder was initialized either with the baseline pretrained encoder weights (Whisper without fine-tuning) or with the encoder weights of the models fine-tuned for verbatim transcription of healthy and patient speech. The ft-h and ft-p model weights were used when predicting accuracy scores on SONIVA-Naming healthy and SONIVA-Naming patient data, respectively, to capture the most relevant speech representations of each group. This setup resulted in a total of 32 models: models were trained separately for each accuracy metric (Semantic, Dysfluency, Self-correction, Phonology), with two encoder weight configurations (fine-tuned or baseline), each in two sizes (Small and Medium) for each dataset separately (SONIVA-Naming healthy and SONIVA-Naming patient). During training, the encoder weights were frozen and only the linear classification head was trained (i.e., linear probing).
Models were trained for a maximum of 8000 steps and with a batch size of 16. A learning rate of $1\times 10^{-5}$ was used, with a linear learning rate scheduler incorporating 500 warm up steps. Models were evaluated on the validation set at 100 step intervals. The model parameters were updated with AdamW optimizer using Cross-Entropy loss function (1). The best model was chosen based on F1 Macro metric on the validation set and was used for final evaluation on the test set.
III-E Evaluation metrics

### v2

# The Second Workshop on GenAI for Health: Potential, Trust, and Policy Compliance

@@ Abstract
Detailed assessment of language impairment following stroke remains a cognitively complex and clinician-intensive task, limiting timely and scalable diagnosis. Automatic Speech Recognition (ASR) foundation models offer a promising pathway to augment human evaluation, but their effectiveness in the context of speech and language impairment remains uncertain. In this study, we evaluate whether Whisper, a state-of-the-art ASR foundation model, can be applied to transcribe and analyze speech from patients with stroke during a picture-naming task. We assess both verbatim transcription accuracy and the model’s ability to support downstream prediction of language function, which has major implications for outcomes after stroke. Our results show that the baseline Whisper model performs poorly on single-word speech utterances. Nevertheless, fine-tuning Whisper significantly improves transcription accuracy (reducing Word Error Rate by 87.72% in healthy speech and 71.22% in speech from patients). Further, learned representations from the model enable accurate prediction of speech quality (average F1 Macro of 0.74 for healthy, 0.75 for patients). However, evaluations on an unseen (TORGO) dataset reveal limited generalizability, highlighting the inability of Whisper to perform zero-shot transcription of single-word utterances on out-of-domain clinical speech and emphasizing the need to adapt models to specific clinical populations. While challenges remain in cross-domain generalization, these findings highlight the potential of foundation models, when appropriately fine-tuned, to advance automated speech assessment and rehabilitation for stroke-related impairments.
1 Introduction
@@
Stroke is one of the leading causes of adult death and disability worldwide, with global incidence of 12.2 million per year [1]. One of the most debilitating and common consequences after stroke is aphasia, an acquired disorder of speech and language production and comprehension, with over 30% prevalence in patients with stroke [2, 3, 4]. Due to high heterogeneity of symptoms, diagnosis of aphasia requires an in-depth assessment to uncover specific impairments and tailor speech therapy to individual patient’s needs. Computerized assessments provide a cost-effective and accessible platform for post-stroke speech evaluation [5], but they often depend on manual transcription and scoring by trained clinicians, which is time-consuming, prone to inconsistencies, and difficult to scale.
@@
Advancements in automatic speech recognition (ASR) methods can fill this gap by automating the speech evaluation pipelines. The extensive linguistic knowledge embedded in foundation ASR models, acquired through pretraining on large and diverse datasets, has the potential to be leveraged in a zero and few shot learning context to automate post-stroke speech evaluation pipelines [6, 7]. However, whilst the latest ASR models perform exceedingly well on healthy speech data [8], several limitations have hindered the translation of ASR to pathological speech analysis. Firstly, current ASR models have been trained on healthy speech, limiting their generalizability to pathological speech. Secondly, the potential variability in speech impairments in aphasia necessitates large clinical speech databases for effective training of ASR models, which are currently relatively limited.
@@
In this study, we aimed to assess the performance of Whisper [8], a state-of-the-art transformer-based ASR model pretrained on 680,000 hours of multilingual data, on a stroke-specific speech database derived from a commonly used Naming task. We fine-tuned Whisper on speech from age-matched healthy older adults and patients with stroke and compared the verbatim transcription performance of the fine-tuned and the baseline Whisper models. Further, we evaluated whether Whisper-derived representations could be effectively applied to clinically-relevant downstream tasks, specifically speech impairment severity prediction. Lastly, we assessed the quality of these predictions by conducting a divergent and convergent validity analysis with patients’ known clinical features.
2 Related Work
@@
Current approaches for automated speech impairment analysis involve three key steps. First, speech-to-text ASR models are employed to generate speech transcriptions from audio recordings. Then, a feature extraction process is applied to the produced transcriptions and audio data. The extracted lexico-acoustic features are used in downstream analyses, such as classification and regression, to estimate the degree of speech impairment [9]. Different algorithms have been used for each step.
2.1 Recurrent Neural Networks
@@
Previous studies have trained customized ASR models to transcribe aphasic speech, subsequently deriving acoustic and transcription-based features to estimate aphasia severity [10, 9]. Le and colleagues [10] utilized a Bidirectional Long-Short Memory-Recurrent Neural Network (BLSTM-RNN) model trained on acoustic features of aphasic speech derived from the AphasiaBank dataset [11] and frame-level senone and monophone labels, obtained through a Hidden Markov Model-Gaussian Mixture Model (HMM-GMM) system, to produce speech transcriptions. Similarly, Qin and colleagues [9] utilized a Time-Delay Neural Network combined with a Bidirectional Long-Short Term Memory (TDNN-BLSTM) model to transcribe Cantonese aphasic speech. The model was trained on acoustic features derived from healthy speech and HMM-GMM-derived triphone target labels [9]. The ASR transcriptions were used to extract features based on text alone [10] or in combination with audio features [9] for downstream estimation of aphasia severity by regression. While the above studies reached promising accuracy, the complex architectures of employed ASR systems required a multi-step training process, limiting global optimization [12]. Moreover, the models were based on RNN variants, which are less effective than Transformers in modeling long-range dependencies [13].
2.2 Transformers
@@
Transformers have revolutionized the field of ASR by implementing self-attention mechanisms, which capture temporal relationships across the entire input sequence simultaneously, thus overcoming the range limitations of RNN-based architectures and enabling more efficient encoding of contextual information [14, 15]. Multiple ASR transformer-based models are currently available for speech-to-text transcription including Transformer Transducer [16], wav2vec 2.0 [17], SpeechStew [18], Conformer [19] and Whisper [8]. However, while transformer-based models demonstrate exceptional performance on healthy speech ASR [8], they often encounter challenges adapting to aphasic speech recognition due to the high heterogeneity in lexico-acoustic features of post-stroke speech. Recent studies have leveraged pretrained transformer-based ASR models, adapting them to enhance performance on aphasic speech. Within this framework[20], XLSR-53, a model built upon wav2vec 2.0 architecture, was fine-tuned to adapt it to aphasic speech transcription. XLSR-53 [21], pretrained on the extensive dataset which contains 56,000 hours of unlabeled audio, uses convolutional neural network layers to generate latent speech representations [17]. These representations are then passed through a transformer network to encode contextualized speech features [17]. However, wav2vec 2.0 architecture is outperformed by other transformer-based foundation models such as Whisper [8, 22]. Whisper-derived representations have been successfully utilized in previous studies to classify dysfluencies in stuttering [23] and to estimate dysarthria severity [24]. Further, Whisper and XLSR-53 model transcripts have been successfully used for relevant speech feature extraction, later used for aphasia type classification, reaching an average F1 score of 90.6 [25]. Recent work also investigated hyperparameter tuning and fine-tuning of Whisper to better fit aphasic speech, reaching promising results of 38.5% and 21.93% WER on aphasic speech transcription, respectively [26, 27]. However, the vast majority of previous work has focused on sentential speech, and the applications of Whisper to single-word utterances remain to be studied.
3 Method
3.1 Dataset
[…]
@@ 3.1 Dataset
An unseen dataset was used to assess model generalizability for single-word pathological speech transcription. This dataset was derived from the TORGO database [31], which includes speech from individuals with dysarthria, a motor speech disorder common in cerebral palsy and amyotrophic lateral sclerosis. Single-word recordings were used to evaluate model transcription of short-form data. The TORGO dataset comprised 1,240 utterances from 7 participants (4 male, 3 female).
3.2 Model
OpenAI’s Whisper [8] was fine-tuned for the verbatim transcription task. For the accuracy classification task, we employed WhisperForAudioClassification, an adaptation that retains the original Whisper encoder but replaces the decoder with a linear classification head, implemented via the Hugging Face Transformers library [32]. Whisper architecture is described in Appendix B.
3.3 Data Pre-processing
@@
All audio data were resampled to 16 kHz. In the main dataset used for model training and evaluation (SONIVA-Naming), recordings without at least one ground truth accuracy metric (Semantic, Dysfluency, Self-correction, Phonology) were excluded when fine-tuning models for the accuracy score prediction task (n = 23 and n = 114 for SONIVA-Naming healthy and patient datasets, respectively). Data were divided into training, validation, and test sets in a 7:1:2 ratio. Splitting was performed at the participant level such that all trials from the same participants were in the same partition. Due to high class imbalance across accuracy scores, the training sets were randomly downsampled to the minority class in each metric. This resulted in 144, 204, 90, and 21 audio files for the SONIVA-Naming healthy dataset and 198, 729, 108, and 171 trials for the patient dataset, respectively.
3.4 Model Training
@@
Whisper was separately fine-tuned on the training set of synthetic, SONIVA-Naming healthy, and SONIVA-Naming patient data, as well as combined datasets of all participant data (SONIVA-Naming healthy and SONIVA-Naming patient) and all available data (synthetic, SONIVA-Naming healthy and SONIVA-Naming patient) (See Appendix C). Small and Medium Whisper models [33, 34] were used to determine the optimal model size. In total, ten fine-tuned models were obtained: fine-tuned on synthetic data (ft-syn), on SONIVA-Naming healthy data (ft-h), on SONIVA-Naming patient data (ft-p), on all participant data (ft-hp) and all available data (ft-all), each in Small and Medium sizes.
Models were trained on a single NVIDIA RTX 6000 GPU for a maximum of 1000 steps, with a batch size of 16, which took approximately 2-4 hours. A learning rate of $1\times 10^{-5}$ was used, with a linear learning rate scheduler incorporating 250 warm-up steps. The model parameters were updated with AdamW optimizer using Cross-Entropy loss function.
Models were evaluated on the validation set (batch size of 8) at 50 step intervals. The best model was chosen based on Word Error Rate (WER) on the validation set and was used for final evaluation on the test sets. The evaluation was completed on four different test sets, namely the synthetic, SONIVA-Naming healthy, and SONIVA-Naming patient as well as the unseen TORGO database.
@@
The setup was formulated as a multi-class classification task, where one of three possible accuracy scores (0, 1 or 2) was predicted for each given Naming task trial recording in the SONIVA-Naming healthy and patient datasets (See Appendix C).
The encoder was initialized either with the baseline pretrained encoder weights (Whisper without fine-tuning) or with the encoder weights of the models fine-tuned for verbatim transcription of healthy and patient speech. The ft-h and ft-p model weights were used when predicting accuracy scores on healthy and patient data, respectively, to capture the most relevant speech representations of each group. This setup resulted in a total of 32 models: models were trained separately for each accuracy metric (Semantic, Dysfluency, Self-correction, Phonology), with two encoder weight configurations (fine-tuned or baseline), each in two sizes (Small and Medium) [33, 34] for each dataset separately (SONIVA-Naming healthy and SONIVA-Naming patient). During training, the encoder weights were frozen and only the linear classification head was trained (i.e., linear probing).
Models were trained on a single NVIDIA RTX 6000 GPU for a maximum of 8000 steps and with a batch size of 16, which took approximately 6-8 hours. A learning rate of $1\times 10^{-5}$ was used, with a linear learning rate scheduler incorporating 500 warm-up steps. Models were evaluated on the validation set at 100 step intervals. The model parameters were updated with AdamW optimizer using Cross-Entropy loss function. The best model was chosen based on F1 Macro metric on the validation set and was used for final evaluation on the test set.
3.5 Evaluation metrics